Theorems of addition of probability and theorems of multiplication of probabilities are used for solving
1. What is the probability of drawing either a king or a queen in a single drawing from a deck of 52 playing cards?
2. There are 3 children in family. The birth o a girl and a boy is considered to be
equally possible. Find out the probability of the fact that all children in the family
are boys.
1
Expert's answer
2020-11-04T16:05:05-0500

1) P(drawing either a king or a queen) = P(drawing a king) + P(drawing a queen) =

"=\frac{4}{52}+\frac{4}{52}=\frac{8}{52}=\frac{2}{13}=0.15"


2) P(3 children are boys):

P(1st child is a boy) * P(2nd child is a boy) * P(3rd child is a boy)="\frac{1}{2}\cdot\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{8}=0.125"